OpenStudy (anonymous):

what's tan10-tan50-tan70 without using trigonometric table

OpenStudy (anonymous):

sqrt(3)

OpenStudy (anonymous):

tan (60+10) ,tan(60-10) looks tan (a+b),tan(a-b)

OpenStudy (anonymous):

tan(a+b)=tana+tan b/1-tan a tan b

OpenStudy (anonymous):

tan(a-b)=tan a-tan b/1+tan atan b

Parth (parthkohli):

It may also be useful to know both of these identities:\[\tan(x) + \tan(60+x)+\tan(120+x) = 3\tan(3x)\]\[\tan(60-x)\tan(x)\tan(60 + x) = \tan(3x)\]